how can i do these problems 8x+6-9x=2-x-15 ,AND 3m-10=2(4M-5),AND 5-(n-4)=3(n+2),and 8(x+1)=4x-8
1
Expert's answer
2012-08-21T07:28:06-0400
Let's consider the first one (all the rest of your tasks are solved similarly):
8x+6-9x=2-x-15
Let's gather terms with x in the left part and the rest in the right
8x+6-9x & =& 2-x-15
& -6 & +x +x -6
hence we obtain:
8x-9x+x=2-15-6
0=-19
The equality we've obtained is false so this problem has no solution.

Another one
3m-10=2(4M-5)
3m-10=8m-10
5m=0 whence m=0

5-(n-4)=3(n+2)
5-n+4=3n+6
3=4n
whence n=3/4